OpenStudy (anonymous):

Consider the piecewise defined function f(x)=x^2+4/x+2 if x<-2 0 if x=-2 2x+3 if x>-2 Calculate f(-3)-f(-2)

OpenStudy (mertsj):

f(-3)=(-3)^2+4/-3+2=13/-1=-13 f(-2)=0 F(-3)-f(-2)=-13-0=-13

OpenStudy (ash2326):

f(-3)-f(-2) for x<-2 f(x)=x^2+4/x+2 so \[f(-3)=\frac{{(-3)}^2+4}{-3+2}\] \[f(-3)=\frac{13}{-1}\] \[f(-3)=-13\] for x=-2 f(x)=0 so f(-2)=0 f(-3)-f(-2)=-13-0=-13

OpenStudy (anonymous):

how did you guys get -13 i keep getting -5....

OpenStudy (mertsj):

What did you get for f(-3) ?

OpenStudy (anonymous):

-5...

OpenStudy (mertsj):

That's your trouble. What is (-3)^2?

OpenStudy (anonymous):

lol ohhh

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw